The average​ profit/loss of a company​ (in thousands of​ dollars) was documented for 30 months. The 5th month showed a loss of ​$3000​, however the 25th month showed a profit of ​$9000. Assuming the​ profit/loss is​ linear, write the equation of the​ line, in​ slope-intercept form, that represents the​ profit/loss for each month x. Use this result to predict the​ profit/loss after 60 months.
​Hint: Use the ordered pairs (5,-3)(25, 9).

Answer by Boreal(15235)

slope is (9-(-3))/(25-5)
12/20=0.6
point slope formula y-y1=m(x-x1), m slope (x1, y1) point
y-9=0.6(x-25)
y=0.6x-15+9
y=0.6x-6, x months and y profit.
for 60 months, y=0.6*60-6 or 36-6=$30000 profit